A bird travels at 500 mph and leaves 10 minutes before his bird friend, which travels at 1500 mphs. How long does it take the friend bird to catch up to the bird that left 10 minutes before him? ... ANSWER NOW IM CURRENTLY IN MY MATH CLASS AND IN BIG TROUBLE IF YOU DONT ANSWER THIS!!
Answer by mananth(16946)

Bird I 500 mph
Bird II 1500 mph
In 10 minutes = 1/6 hour bird I travels 500*1/6 = 83.33 miles.
difference between their speeds = 1000
d= 83.33
speed = 1000 mph
time = d/r
83.33/1000
0.0833 hours
0.0833*3600= 300 seconds
time to catchup = 300 seconds